OpenStudy (anonymous):

What day of the week was this day July 4th 1776?

OpenStudy (anonymous):

July 4, 1776 is the 186th day of the year 1776 in the Gregorian calendar. There are 180 days remaining until the end of this year. The day of the week is Thursday.

OpenStudy (anonymous):

U want to calculate it

OpenStudy (lgbasallote):

Independence Day of U.S.

OpenStudy (anonymous):

I need to know how to solve it

OpenStudy (anonymous):

U need to write a program

OpenStudy (anonymous):

I think Once I did this type of question for @yahoo

OpenStudy (anonymous):

here is an example of one they did in class i cant figure out how to solve it for this one though July 6,1994 #of complete years 1993 same as 393(1993 mod400=393) 300 yrs have 1 odd day 93yrs have (93+(93/4)) mod 7 odd days =93+23mod7 odd days=4 thursday

OpenStudy (anonymous):

It's Thursday

OpenStudy (anonymous):

can you tell me how you figured it out?

OpenStudy (anonymous):

http://wiki.answers.com/Q/What_day_of_the_week_was_July_4_1776 That's how. Lol